The Structure of Our World

I look at our world as having 3 distinct distance scales

1. Very largest scales • Size of the known universe (∼ 1025 m)

• Galactic clusters (∼ 1022 m)

• Galaxies (∼ 1021 m)

• Solar System (∼ 1013 m)

2. “Mushy” intermediate scales • The Earth (∼ 107 m)

• You and me (∼ 1 m)

• Virus (∼ 10−7 m)

3. Very smallest scales • Molecules (∼ 10−9 m)

• Atom (∼ 10−10 m)

• Nucleus of an atom (∼ 10−14 m)

Big differences in the natural structures

• But very largest and smallest scales very similar – “simple” structures

– have very related problems

The Cosmic Contradiction

The Big Bang happened –

• just energy (matter and anti-matter) to begin with • “hot soup” rapidly expanded and cooled

• protons and neutrons formed in 1 s

• atoms started forming after about 300,000 years – radiation and matter stopped interacting – radiation continued cooling undisturbed – matter then started forming structure

Radiation is still kicking around today

• Cosmic microwave background discovered in 1965

• Has a temperature of about 3∘ K (−270∘ C)

• Turns out to be VERY smooth

Implies Universe was also smooth at 300,000 years

• Can’t explain the large structures we see

• Need at least small “ripples” that then grow

Why are Heavy Quarks Hard to Find?

Hard to produce

• Must sift through lots of matter-antimatter collisions

• Happens about 1 in every billion collisions?

They Decay!

• A top quark decays via the weak force into a bottom quark • The bottom quark then decays into lighter quarks

Ya can’t see free quarks

• Quarks are “not allowed” to appear unescorted in public • Have to “dress” themselves up around other quarks

• Result in “jets” of particles

Summary

Haven’t reached the end of the story

• Many open questions, for example – What and where is mass?

– How did our universe become what it is today?

– What really are quarks?

But we are making progress

• We have a “picture” that works pretty well already • Focusing on those pieces that don’t quit fit

Soon we will have a much better understanding of

• the very big,

• the very small,

• ...and all that we find in between!